find the slope, and y-intercept on the graph, of the line 5x-3y=6
Answer by Quadratic1600(28)

Hi
Right so for this graph line we have to rearrange it in to the form y = mx + c or in some cases it is written as
y = mx + b (It doesn't matter here)
So 5x - 3y = 6 equals
3y = 5x - 6
Now divide by 3 to get
y = 5/3x - 2
So the gradient/slope is 5/3 and the y intercept is -2